OpenStudy (anonymous):

Determine an equation of the line through the given point is parallel to the given line: (-1,2): 3y+2x=6

OpenStudy (anonymous):

What's the slope of the given line?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

Can you quickly solve for y?

OpenStudy (anonymous):

3y=6+2x

OpenStudy (anonymous):

If you put the equation in the form y=mx + b m will be your slope.

OpenStudy (anonymous):

err y=mx+b m is the slope.

OpenStudy (anonymous):

That's not quite right. It should be 3y=-2x + 6

OpenStudy (anonymous):

Now divide by 3.

OpenStudy (anonymous):

So y=-2x+2

OpenStudy (anonymous):

So y=-2x+2

OpenStudy (anonymous):

No. \[y=\frac{-2}{3}x + 2\]

OpenStudy (anonymous):

Go a little slower, you 're making a number of clerical mistakes.

OpenStudy (anonymous):

Okay, I understand. Now, isn't there yet another formula to put in the points?

OpenStudy (anonymous):

Once you have the slope of the given line you know the slope of your parallel line (parallel lines have the same slope). You can make your new line using the point slope formula: \[y-y_0 = m(x-x_0)\] Where m is the slope of your line, and \(x_0,y_0\) are the x and y values of a point on the line.

OpenStudy (anonymous):

So m=the -2/3+2

OpenStudy (anonymous):

Sorry, i'm horrible at this

OpenStudy (anonymous):

y-2=m(x-(-1))

OpenStudy (anonymous):

Then y-2=m(x+1)

OpenStudy (anonymous):

m = -2/3 Not the +2.

OpenStudy (anonymous):

-2/3 is the slope. y = mx + b The m is the slope.

OpenStudy (anonymous):

So is this right then.....y=-2/3s11

OpenStudy (anonymous):

y-2=-2/3(x+1) is correct. Then simplify. into y = something.

OpenStudy (anonymous):

Wow, i don't know how else to simplify this

OpenStudy (anonymous):

\[y-2 = \frac{-2}{3}(x+1)\] Add 2 to both sides. \[y = \frac{-2}{3}(x+1) + 2 \] Multiply through the -2/3. \[y = \frac{-2}{3}x -\frac{2}{3} + 2 \] Combine the non-variable terms.

OpenStudy (anonymous):

So y=2?

OpenStudy (anonymous):

Okay, i'm seriously horrible at this.

OpenStudy (anonymous):

Is there a Algebra for Dummies book out there? I'm sure there is, and it's going to be with my other books.

OpenStudy (anonymous):

There's a number of good videos you can watch.

OpenStudy (anonymous):

Start there and work on some of the more complicated equations too

OpenStudy (anonymous):

Well, I need it seriously to finish this class. . Can you help me finish this one please.....

OpenStudy (anonymous):

I will certainly do that.

OpenStudy (anonymous):

Just bookmarked it.

OpenStudy (anonymous):

You should have y = (-2/3)x + 4/3

OpenStudy (anonymous):

All the other questions are coming up similar to 3x+y-6=0 They all have 0's at the ends..

OpenStudy (anonymous):

That doesn't matter. You can always move things back and forth across the equal sign by adding or subtracting things.

OpenStudy (anonymous):

So I could do -2/3x+y+2=0

OpenStudy (anonymous):

You can, but there's no need.

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

Also that would be wrong.

OpenStudy (anonymous):

Because you have to add (2/3)x -4/3 to both sides (or subtract y from both sides) to get 0.

OpenStudy (anonymous):

Okay, i understand that. Thanks.

OpenStudy (anonymous):

You sure are good at this.

OpenStudy (anonymous):

You need to review working with equations a bit, and take things slowly, but you'll get there. Math is a practiced skill, not an innate talent.

OpenStudy (anonymous):

I know, thanks again.